OpenStudy (anonymous):

Could you help please? I need to evaluate the following integral of cos sq root of x/sq root dx Thank you

OpenStudy (anonymous):

\[u=\sqrt{x}, du=\frac{1}{2\sqrt{x}}dx, 2du=\frac{dx}{\sqrt{x}}\] should do it

OpenStudy (raffle_snaffle):

 satellite when he can't figure the problem out

OpenStudy (anonymous):

I'm sorry that is a cosine in front of \[\sqrt{x}\]

OpenStudy (anonymous):

i know i assumed it was \[\large \int \frac{\cos(\sqrt{x})}{\sqrt{x}}dx\] right?